OpenStudy (anonymous):

The population (in millions) of a certain island can be approximated by the function: P(x)=50 (1.05)^x Where x is the number of years since 2000. In which year will the population reach 200 million?

OpenStudy (anonymous):

it would be 2028, right? or am i off?

OpenStudy (wolf1728):

population = 50*(1.05)^x population / 50 = (1.05)^x log (population / 50) = x * log (1.05) log (population / 50) / log (1.05) = x we need to know when population = 200,000,000 so according to the formula the population will equal 200: log (200 / 50) / log (1.05) = x 0.6020599913 / 0.0211892991 = x x = 28.4133981658 Years